Status Report

NASA Space Station On-Orbit Status 7 Nov 2003

By SpaceRef Editor
November 7, 2003
Filed under , , ,
NASA Space Station On-Orbit Status 7 Nov 2003
iss

All ISS systems continue to function nominally, except as noted previously or below. Our Russian partners are observing a Holiday (“Day of Consent and Reconciliation”).

At the SVO-ZV water supply system in the Russian segment (RS), FE Alexander Kaleri completed the three-day activity of washing and sterilizing SVO-ZV water supply equipment, using disinfectant delivered on 12P. He closed off by collecting a water sample from the SVO-ZV for the subsequent microbial analysis by U.S. samplers. [When Kaleri replaced the SVO-ZV receptacle tip, he observed a small leak in it, which disappeared when he took the micro sample (could be due to particulate/contamination on the nozzle valve, to be taken into account during ground processing of the sample).]

CDR Michael Foale undertook the monthly water sampling for in-flight chemistry/microbiology analysis, using approved Russian sampling procedures with the U.S. WS&A (water sampler & archiver) for collection and the WMK (water microbiology kit) for treatment within 6 hours of the collection. Results will be available after a two-day incubation period. [Samples were taken in the SM at the potable water SRV-K hot port and from the EDV container of the SVO-ZV water supply system (see above). Last time done: 9/24.]

After the water collection activity, the crew reclaimed the water from the sample and waste bags by using an absorbent item (such as a used T-shirt) to soak it up and squeeze it into a bag for “reuse” as condensate water. This conserves the onboard water and prevents stowage issues inside the already full WMK and WS&A kits.

Kaleri cycled Russian laptop batteries for maintenance, with discharge and recharge required to be scheduled six hours apart.

Foale performed the regular once-a-week maintenance reboot on the operational PCS (portable computer system) laptops in the station.

Mike also conducted the periodic inventory audit of the available CWCs (contingency water containers) and their contents, to keep track of onboard water supplies.

Today marked the start of the NASA/JSC renal (kidney) stone experiment activities by the Expedition 8 crew. In preparation of the new round, the crew first reviewed the Renal CBT (computer-based training). Michael Foale then unstowed and set up the experiment hardware. Both he and Sasha Kaleri began their first day of the investigation by starting their diet logs.   [This long-range preventive medicine investigation features daily random ingestion of either potassium citrate or placebo tablets. It is Dr. Peggy Whitson’s double-blind research study investigating methods to prevent formation of kidney stones in zero-G. Part of the experiment consists in keeping a metabolic diet log (food and fluid intake), followed by collection of urine samples several times per day during each week-long session.]

FE Kaleri performed another session of the Russian Uragan earth observations program, using the LIV Betacam video system from SM window #9. [The window is always protected by its closed internal cover, except for Earth observation periods. Both internal and external covers were closed after the conclusion of today’s session. Targets of interest today were the Galapagos Islands, the coastal line of South America and the Andes, panoramic views of the mountainous country of the Andes, landscape mapping of South America, and the valley of the Parana River.]

Mike Foale shut down the EarthKAM experiment for today and closed the Lab window shutter.

Sasha conducted the daily routine maintenance of the SM’s SOZh life support system (including ASU toilet facilities) and prepared the IMS (inventory management system) “delta” list for automatic export/import to update the database, while Mike conducted the regular routine status checkup of the autonomous Lab payloads (currently only PCG-STES010).

Foale called down the daily “ad hoc” cabin air ppO2 (oxygen partial pressure) measurement of the CSA-CP (compound specific analyzer-combustion products), used by the ground for trending analyses, followed by the weekly calldown of the ppCO2 value, measured by the CDMK (carbon dioxide monitoring kit).

Mike transferred data files from the physical exercise equipment (TVIS and RED) to the MEC (medical equipment computer). Subsequently, he also performed the periodic transfer of accumulated data files from the CMS HRM (crew medical systems/heart rate monitor) to the MEC (medical equipment computer), then deleted them on the HRM.

Sasha conducted the regular 20-min. tag-up with ground specialists via S-band to discuss the weekly IMS (inventory management system) status update. [Point of discussion was the current locations of Soyuz TMA-3 payloads shown in the IMS database, viz., five containers with DVDs, a PDA with cable, player and headset, a Relaksatsiya and Reflotron-4 kit, personal items, etc.]

At about 2:10pm EST, the crew will have their weekly teleconference with the ISS Flight Director at MCC-H.

Both crewmembers completed their daily physical exercise program (2.5 hrs.) on TVIS treadmill, CEVIS cycle ergometer, and RED anaerobic exerciser.

Due to a procedural error, yesterday’s charging of defibrillator battery #1007 could not be completed. The procedure will be changed (to include a crew calldown at Charge + 1 hour) and used for the next battery charge event.

During yesterday’s videotape playback on VTR 2 (video tape recorder #2), the picture became “jittery”, after being OK earlier. VTR2 was placed in Standby and cannot be used (losing VTR redundancy) until the problem is resolved.

At 8:21pm EST tonight, the station, currently continuing in earth-fixed LVLH flight attitude, will maneuver again to sun-oriented XPOP (x-axis perpendicular to orbit plane).

Due to the reduction in flow of solar protons of >10 MeV energy to fewer than 10 particles per second per square meter, TsUP/Moscow cancelled the radiation alert this morning, terminating the need for additional readings with the two “Pille” was radiation dosimeters. [They were to be returned to their original locations (in the ASU toilet and on the SM control panel).]

Today’s CEO (crew earth observation) targets, taking into account the current LVLH attitude, and including the targets of the Lewis & Clark 200-year memorial locations, wereEastern Mediterranean Aerosol (high pressure is setting up over central Europe, so that smog should build up in the Po River valley of northern Italy. Looking left from the crew’s position over Rome to the Adriatic Sea on the far side of Italy’s boot. The oblique look angle and low sun angle both assisted in imaging smog masses. Slight overexposure also helps. Po Valley industrial smog commonly flows as visible masses into the Eastern Mediterranean, with transport documented north into the Black Sea, and also east to Lebanon and even into Iraq),SE Africa Aerosol(high atmospheric pressure conditions are established over southern Africa. Looking left towards the Witwatersrand complex of cities. Severe smog appears in southern Africa as a purplish mass, and is a mixture of smoke and dust. The masses can reach subcontinental size. Obliques left over the Lesotho Mountains and coastal lowlands should have shown significantly clearer air),Recife, Brazil(looking a touch left on the coast),Guadalajara, Mexico(looking a touch left, for this city located on the edge of a major canyon),Buenos Aires, Argentina(nadir pass),Kingman Reef, Central Pacific(looking slightly left for detailed images of this reef, needed for the global reef mapping project. This US reef encloses a sheltered lagoon, which served as a way station for flying boats on Hawaii-American Samoa run in the 1930s. The reef is often awash [and thus a designated maritime hazard], and no land plants grow here, but it does support abundant marine fauna and flora. In 2001, the waters were designated a US National Wildlife Refuge),Palmyra Atoll, Central Pacific(looking slightly left for images of this reef, also part of the global reef mapping project. This US atoll measures 12 sq.km in area and is completely surrounded by a coral reef, which was designated a National Wildlife Refuge in 2001), andJarvis Island, Central Pacific(looking slightly right for detailed images of this reef, needed for the global reef mapping project. This 4.5-square-km atoll provided tons of guano in the 1800s. It is US territory now designated a National Wildlife Refuge. The highest point is only 6 feet above mean sea level).

CEO images can be viewed at the websites.

See also the website “Space Station Challenge” at

U.S. and Russian Segment Status (as of 3:00am EST).

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

  • Elektron O2 generator is powered On, 18A. Vozdukh CO2 scrubber is On (in Manual Mode 5/3). U.S. CDRA CO2 scrubber is on Standby (ready in dual-bed mode). TCCS (trace contaminant control subsystem) is operating. MCA (major constituents analyzer) is off (in Life Extending Mode). BMP Harmful Impurities unit: absorbent bed #1 in Purify mode, bed #2 in Purify mode. RS air conditioner SKV-1 is On; SKV-2 is Off.
  • SM Working Compartment:  Pressure (mmHg) — 742; temperature (deg C) — 27.2; ppO2 (mmHg) — 159.5; ppCO2 (mmHg) — 2.5.
  • SM Transfer Compartment:  Pressure (mmHg) — 761; temperature (deg C)  20.4.
  • FGB Cabin:  Pressure (mmHg) — 756; temperature (deg C) — 23.0.
  • Node:  Pressure (mmHg) — 747.07; temperature (deg C) — 22.9 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • U.S. Lab:  Pressure (mmHg) — 749.06; temperature (deg C) — 25.0; ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • Joint Airlock (Equip. Lock):   Pressure (mmHg) — 750.65; temperature (deg C) — 23.6; shell heater temp (deg C) — 21.6, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • PMA-1:  Shell heater temp (deg C) — 22.5
  • PMA-2:  Shell heater temp (deg C) — 20.7

(n/a = data not available)

Electrical Power Systems (EPS):

  • Both P6 channels fully operational. BGA (beta gimbal assembly) 2B and 4B both in Directed (“triangle mode”) Position (non-suntracking). 
  • SM batteries: (data unavailable).
  • FGB batteries:  Batteries #1 is off; all other batteries (5) are in “Partial Charge” mode.
  • Plasma Contactor Unit PCU-1 in Standby mode; PCU-2 in Standby mode.

Command & Data Handling Systems:

  • C&C-2 MDM is prime, C&C-1 is back-up, and C&C-3 is in standby.
  • GNC-1 MDM is prime; GNC-2 is Backup.
  • INT-2 is operating; INT-1 is Off.
  • EXT-2 is On (primary), EXT-1 is Off (both now upgraded to R3).
  • LA-1, LA-2 and LA-3 MDMs are all operating.
  • PL-1 MDM is Off; PL-2 MDM is Operational.
  • APS-1 (automated payload switch #1) and APS-2 are both On.
  • SM Terminal Computer (TVM): 3 redundant lanes (of 3) operational (all lanes reintegrated 11/5).
  • SM Central Computer (TsVM): 2 redundant lanes (of 3) operational; string #3 dropped out 10/22.

Propulsion System:

  • Total propellant load available:3734 kg (8232 lb) as of 11/6 [SM(755) + FGB(2627) + Progress M(352) + Progress M-1(0)]. (Capability: SM — 860 kg; FGB — 6120 kg).

Attitude Control Systems:

  • 3 CMGs on-line (CMG-1 failed).
  • State vector source — U.S. SIGI-1 (GPS)
  • Attitude source — U.S. SIGI-1 (GPS)
  • Angular rate source — RGA-1

Flight Attitude:

  • LVLH YVV (y-axis in velocity vector, i.e. flying “sidewise”) [yaw: -90 deg, pitch: -8.9 deg., roll: 1.8 deg]), with CMG Momentum Management.

Communications & Tracking Systems:

  • FGB MDM-1 is powered Off; FGB MDM-2 is operational.
  • All other Russian communications & tracking systems are nominal.
  • S-band is operating nominally (on string 2).
  • Ku-band is operating nominally.
  • Audio subsystem is operating nominally (IAC-2 is prime, IAC-1 is suspect).
  • Video subsystem operating nominally.
  • HCOR (high-rate communications outage recorder) is operating nominally.

Robotics:

  • SSRMS/Canadarm2 based at MBS PDGF #2/LEE A, with Keep Alive (KA) power on both strings.
  • MBS: KA power on both strings. 
  • MT: latched and mated at WS4.
  • POA: KA power on both strings.
  • RWS (robotics workstations): Lab RWS is On (DCP connected); Cupola RWS is Off.

ISS Orbit  (as of last night, 11:37pm EST [= epoch]):

  • Mean altitude — 376.8 km
  • Apogee — 381.1 km
  • Perigee — 372.4 km
  • Period — 92.1 min.
  • Inclination (to Equator) —  51.63 deg
  • Eccentricity — 0.000645
  • Solar Beta angle — 8.2 deg (magnitude increasing)
  • Orbits per 24-hr. day — 15.64
  • Mean altitude loss in last 24 hours — 190 m
  • Revolutions since FGB/Zarya launch (Nov. ’98)  — 28335
  • For more on ISS orbit and worldwide naked-eye visibility dates/times, see http://www.hq.nasa.gov/osf/station/viewing/issvis.html

SpaceRef staff editor.